# Hopfield Neural Networks for Online Constrained Parameter Estimation with Time-Varying Dynamics and Disturbances

### v1

@@ I-A Problem setting and motivation
  • PB–RLS (Projection-Based RLS). A Recursive Least-Squares (RLS) update is followed by a projection onto the feasible set to enforce linear constraints on parameters [1, 2, 3, 4, 5]. PB–RLS is computationally light and tracks drifts via forgetting, but it does not explicitly model disturbances and typically assumes linear–Gaussian noise.
  • DA–PB–KF (Disturbance-Augmented, Projection-Based Kalman Filter). A Kalman filter is augmented with bias/unknown input states to absorb additive disturbances, and each update is projected to satisfy parameter constraints (cf. constrained KF frameworks [6, 7]). This improves disturbance rejection over PB–RLS at a moderate extra cost, while retaining sensitivity to modelling/non-Gaussian effects.
  • DA–PB–MHE (Disturbance-Augmented, Projection-Based Moving Horizon Estimation). A finite-horizon optimisation (MHE) with explicit constraints and disturbance/parameter augmentation, standard in MHE formulations, provides a strong accuracy/robustness baseline under constraints, additive disturbances, and slow parameter drift, albeit with the highest online computational burden (see, e.g., [8, 9, 10, 11]).
For completeness, we also reference some other widely used general-purpose estimators. Least Mean Squares (LMS) is computationally frugal with effective frequency-domain variants [12, 13], yet converges slowly and is not inherently constrained [4]. KF/CEKF families achieve Minimum Mean Square Error (MMSE) under Gaussian assumptions and admit constrained projections [6, 7, 14, 15], at higher computational cost and sensitivity to model errors. Particle filters address non-linear/non-Gaussian settings and can encode constraints [16, 17, 18], but often scale poorly in real time [19].
@@ I-B Hopfield neural networks
Since the seminal work of Hopfield and Hopfield–Tank [20, 21], analogue recurrent networks have been used both as associative memories and as neurodynamic optimisers, with numerous applications in the 1990s, e.g., [22, 23, 24, 25, 26]. Later, high-throughput implementations on GPUs/FPGAs [27, 28, 29] were proposed. More recently, theoretical developments reinterpret modern Hopfield layers (dense associative memories) as attention-like mechanisms with higher-order energies [30, 31]. In parallel, differentiable optimisation layers embed convex programmes in deep models [32, 33], and control-orientated works learn Lyapunov functions and stabilising policies [34, 35].
@@
which acts as a Lyapunov function [21, 36].
A standard way to use an HNN for online Least-Squares (LS) with $w=W\theta\;$ is to identify $v=\theta$ and set $T=-W^{\top}W$, $b=W^{\top}w$. Although simple, this mapping is ill-suited for constrained, time-varying operation: (i) $T(t),b(t)$ vary with the data, so (2) may no longer be a Lyapunov function; (ii) $W^{\top}W$ can be ill-conditioned, amplifying noise along weakly excited directions; (iii) equality/inequality constraints are not enforced natively. Prior HNN-based parameter estimators [37, 38, 39, 40, 41, 42, 43, 44] mainly focus on unconstrained problems and do not provide stability guarantees under time variation and disturbances.
@@ II Mapping Online Constrained Estimation with Disturbances onto a Hopfield Neural Network
penalises deviations from the valid subspace $\{x:\,Mx=m\}$ [24]. In what follows, we instantiate this construction with $M=W$ (data subspace) and with $M=A$ (constraint subspace, with inequalities lifted by slack neurons), yielding projector–based Hopfield weights and bias that are updated online from $(W(t),w(t))$ while enforcing (time-invariant) linear constraints natively.
@@ III-A Global bounds
There exists $c_{\star}>0$ such that $c_{\star}\leq\inf_{t}\lambda_{\min}\!\big(P_{W}(t)+\eta P_{A,\theta}\big)$, or equivalently $\operatorname{rank}\!\big[\begin{smallmatrix}W(t)\\ A_{\theta}\end{smallmatrix}\big]=p$, where $p$ is the number of parameters to estimate. The joint identifiability (or constraint–augmented full rank) condition ensures contraction in parameter space (dimension $p$). In fact, for orthogonal projectors $P_{U},P_{V}$ onto subspaces $U,V\subset\mathbb{R}^{p}$, the sum $P_{U}+\eta P_{V}$ ($\eta>0$) is positive definite iff $U+V=\mathbb{R}^{p}$, or equivalently, the stacked matrix of any bases of $U$ and $V$ has full column rank [45, 46, 47, 48] (cf. Lemma 1).
@@ IV Simulation Results
We evaluate the proposed HNN estimators on a 2-DOF mass–spring–damper (MSD) plant (Fig. 1; cf. [49, 50]) with unknown constants $k_{1},b_{1},k_{2},b_{2}$, a disturbance $d(t)\sim\mathcal{N}(\mu,\sigma^{2})$ (white Gaussian), and a slowly time–varying stiffness $k_{1}(t)$ and compare with the performance of PB-RLS, DA-PB-KF and DA-PB-MHE algorithms. Subsection B presents the behaviour of the standard LS (unconstrained) HNN. Subsections C–E compare the proposed HNN estimators against PB–RLS (constraints only), DA–PB–KF (constraints+disturbances), and DA–PB–MHE (constraints+disturbances+time–varying parameters) on a common setup. Subsection F reports a 10–trial Monte Carlo across varying initial conditions, disturbance realizations, and parameter drift. We conclude with a brief complexity comparison of all four methods.
@@ IV-A Dynamical System and Estimation Model
We use the 2-DOF mass–spring–damper (MSD) of Fig. 1 (cf. [49, 50]): two masses $m_{1},m_{2}$ coupled by $(k_{1},b_{1})$, with mass 2 attached to ground by $(k_{2},b_{2})$. The first mass is actuated by a known force $f$, and an additive disturbance $d$ may act on the second mass. Let $x=[x_{1},x_{2},\dot{x}_{1},\dot{x}_{2}]^{\top}$ and $u=[f,d]^{\top}$. The masses $m_{1},m_{2}$ are known; $(k_{1},b_{1},k_{2},b_{2})$ are unknown to be estimated. We consider scenarios with $d(t)=0$, $d(t)\sim\mathcal{N}(\mu,\sigma^{2})$ (constant ++ white Gaussian noise) and a slowly time–varying $k_{1}(t)$.
@@ IV-G Complexity analysis and parallelisation
The per-step work and parallel time are summarised in Table II. PB-RLS follows the classical projected-based RLS literature [51, 52, 1, 2, 3, 4, 5]; DA-PB-KF reflects constrained/augmented Kalman formulations [6, 7, 14, 15]; DA-PB-MHE follows canonical MHE techniques [8, 9, 10, 11].
The per-step work and parallel-time estimates in Table II follow standard operation counts and solver scalings. For all methods, projector/innovation steps are applied via solves: two GEMV and two triangular solves using the Cholesky factors of $WW^{\top}$ (online) and $AA^{\top}$ (offline), rather than forming explicit projectors. The arithmetic costs for Cholesky, triangular solves, and BLAS matrix–vector/matrix–matrix primitives are taken from [53]. The per-sample complexity of RLS that underlies the PB-RLS row follows [54]. The dense Kalman filter innovation, gain, and covariance-update costs that appear in DA-PB-KF come from [55]. For MHE, the fact that each step solves a horizon-$N$ constrained least-squares/QP (dense $\mathcal{O}(N(p{+}q)^{3})$ vs. sparse/structured factorizations) is based on [8, 9] and the MPC text [10]. The real-time iteration viewpoint that justifies the sparse/condensed complexity discussion is due to [56]. Parallel time is reported as an order-of-growth proxy for the critical path using the work–span (Brent) bound,
@@
with $W$ the total work, $D$ the span (depth), and $P$ the number of processors [57].
@@
The proposed HNN estimators are a compelling middle ground: they inherit the parallel efficiency of dense linear–algebra workloads, enforces constraints by construction, mitigates disturbance bias via compensation neurons, and comes with explicit GUUB guarantees. Against the three baselines commonly used in practice [51, 52, 1, 2, 3, 4, 5, 6, 7, 14, 15, 8, 8, 9, 10, 11], it offers a competitive compromise between performance and online complexity, especially when implemented on parallel hardware.

### v2

@@ I-A Problem setting and motivation
  • PB–RLS (Projection-Based RLS). A Recursive Least-Squares (RLS) update is followed by a projection onto the feasible set to enforce linear constraints on parameters [MURAKAMI1998, ZHU1999, HASSAN2009, UHLICH2010, WANG2018]. PB–RLS is computationally light and tracks drifts via forgetting, but it does not explicitly model disturbances and typically assumes linear–Gaussian noise.
  • DA–PB–KF (Disturbance-Augmented, Projection-Based Kalman Filter). A Kalman filter is augmented with bias/unknown input states to absorb additive disturbances, and each update is projected to satisfy parameter constraints (cf. constrained KF frameworks [LUO2012, JOUKOV2015]). This improves disturbance rejection over PB–RLS at a moderate extra cost, while retaining sensitivity to modelling/non-Gaussian effects.
  • DA–PB–MHE (Disturbance-Augmented, Projection-Based Moving Horizon Estimation). A finite-horizon optimisation (MHE) with explicit constraints and disturbance/parameter augmentation, standard in MHE formulations, provides a strong accuracy/robustness baseline under constraints, additive disturbances, and slow parameter drift, albeit with the highest online computational burden (see, e.g., [RAO2001, RAO2003, RAWLINGS2017, ZAVALA2008]).
For completeness, we also reference some other widely used general-purpose estimators. Least Mean Squares (LMS) is computationally frugal with effective frequency-domain variants [ARABLOUEI2015, PENG2024], yet converges slowly and is not inherently constrained [UHLICH2010]. KF/CEKF families achieve Minimum Mean Square Error (MMSE) under Gaussian assumptions and admit constrained projections [LUO2012, JOUKOV2015, HRUSTIC2020, SONG2022], at higher computational cost and sensitivity to model errors. Particle filters address non-linear/non-Gaussian settings and can encode constraints [PAPI2012, DEMIRBAS2015, HU2023], but often scale poorly in real time [SATOH2024].
@@ I-B Hopfield neural networks
Since the seminal work of Hopfield and Hopfield–Tank [HOPFIELD1982, HOPFIELD1985], analogue recurrent networks have been used both as associative memories and as neurodynamic optimisers, with numerous applications in the 1990s, e.g., [TAGLIARINI1991, ALI1993, GEE1994, SMITH1999, ZHANG2014]. Later, high-throughput implementations on GPUs/FPGAs [BASTOSFILHO2011, MEI2014, YAO2014] were proposed. More recently, theoretical developments reinterpret modern Hopfield layers (dense associative memories) as attention-like mechanisms with higher-order energies [KROTOV2016, RAMSAUER2021]. In parallel, differentiable optimisation layers embed convex programmes in deep models [AMOS2017, AGRAWAL2019], and control-orientated works learn Lyapunov functions and stabilising policies [CHANG2019, DAI2021].
@@
which acts as a Lyapunov function [HOPFIELD1985, ABE1989].
A standard way to use an HNN for online Least-Squares (LS) with $w=W\theta\;$ is to identify $v=\theta$ and set $T=-W^{\top}W$, $b=W^{\top}w$. Although simple, this mapping is ill-suited for constrained, time-varying operation: (i) $T(t),b(t)$ vary with the data, so (2) may no longer be a Lyapunov function; (ii) $W^{\top}W$ can be ill-conditioned, amplifying noise along weakly excited directions; (iii) equality/inequality constraints are not enforced natively. Prior HNN-based parameter estimators [ATENCIA2004, ATENCIA2005, HU2005, ALONSO2007, ALONSO2009, ATENCIA2013, ATENCIA2015, BURAKOV2018] mainly focus on unconstrained problems and do not provide stability guarantees under time variation and disturbances.
@@ II Mapping Online Constrained Estimation with Disturbances onto a Hopfield Neural Network
penalises deviations from the valid subspace $\{x:\,Mx=m\}$ [GEE1994]. In what follows, we instantiate this construction with $M=W$ (data subspace) and with $M=A$ (constraint subspace, with inequalities lifted by slack neurons), yielding projector–based Hopfield weights and bias that are updated online from $(W(t),w(t))$ while enforcing (time-invariant) linear constraints natively.
@@ III-A Global bounds
There exists $c_{\star}>0$ such that $c_{\star}\leq\inf_{t}\lambda_{\min}\!\big(P_{W}(t)+\eta P_{A,\theta}\big)$, or equivalently $\operatorname{rank}\!\big[\begin{smallmatrix}W(t)\\ A_{\theta}\end{smallmatrix}\big]=p$, where $p$ is the number of parameters to estimate. The joint identifiability (or constraint–augmented full rank) condition ensures contraction in parameter space (dimension $p$). In fact, for orthogonal projectors $P_{U},P_{V}$ onto subspaces $U,V\subset\mathbb{R}^{p}$, the sum $P_{U}+\eta P_{V}$ ($\eta>0$) is positive definite iff $U+V=\mathbb{R}^{p}$, or equivalently, the stacked matrix of any bases of $U$ and $V$ has full column rank [HornJohnson2013, Bhatia2007, Halmos1969, Ljung1999] (cf. Lemma 1).
@@ IV Simulation Results
We evaluate the proposed HNN estimators on a 2-DOF mass–spring–damper (MSD) plant (Fig. 1; cf. [FEKRI2006, FEKRI2007]) with unknown constants $k_{1},b_{1},k_{2},b_{2}$, a disturbance $d(t)\sim\mathcal{N}(\mu,\sigma^{2})$ (white Gaussian), and a slowly time–varying stiffness $k_{1}(t)$ and compare with the performance of PB-RLS, DA-PB-KF and DA-PB-MHE algorithms. Subsection B presents the behaviour of the standard LS (unconstrained) HNN. Subsections C–E compare the proposed HNN estimators against PB–RLS (constraints only), DA–PB–KF (constraints+disturbances), and DA–PB–MHE (constraints+disturbances+time–varying parameters) on a common setup. Subsection F reports a 10–trial Monte Carlo across varying initial conditions, disturbance realizations, and parameter drift. We conclude with a brief complexity comparison of all four methods.
@@ IV-A Dynamical System and Estimation Model
We use the 2-DOF mass–spring–damper (MSD) of Fig. 1 (cf. [FEKRI2006, FEKRI2007]): two masses $m_{1},m_{2}$ coupled by $(k_{1},b_{1})$, with mass 2 attached to ground by $(k_{2},b_{2})$. The first mass is actuated by a known force $f$, and an additive disturbance $d$ may act on the second mass. Let $x=[x_{1},x_{2},\dot{x}_{1},\dot{x}_{2}]^{\top}$ and $u=[f,d]^{\top}$. The masses $m_{1},m_{2}$ are known; $(k_{1},b_{1},k_{2},b_{2})$ are unknown to be estimated. We consider scenarios with $d(t)=0$, $d(t)\sim\mathcal{N}(\mu,\sigma^{2})$ (constant ++ white Gaussian noise) and a slowly time–varying $k_{1}(t)$.
@@ IV-G Complexity analysis and parallelisation
The per-step work and parallel time are summarised in Table II. PB-RLS follows the classical projected-based RLS literature [ASKARI2016, REICHBACH2016, MURAKAMI1998, ZHU1999, HASSAN2009, UHLICH2010, WANG2018]; DA-PB-KF reflects constrained/augmented Kalman formulations [LUO2012, JOUKOV2015, HRUSTIC2020, SONG2022]; DA-PB-MHE follows canonical MHE techniques [RAO2001, RAO2003, RAWLINGS2017, ZAVALA2008].
The per-step work and parallel-time estimates in Table II follow standard operation counts and solver scalings. For all methods, projector/innovation steps are applied via solves: two GEMV and two triangular solves using the Cholesky factors of $WW^{\top}$ (online) and $AA^{\top}$ (offline), rather than forming explicit projectors. The arithmetic costs for Cholesky, triangular solves, and BLAS matrix–vector/matrix–matrix primitives are taken from [GolubVanLoan2013]. The per-sample complexity of RLS that underlies the PB-RLS row follows [Sayed2003]. The dense Kalman filter innovation, gain, and covariance-update costs that appear in DA-PB-KF come from [Simon2006]. For MHE, the fact that each step solves a horizon-$N$ constrained least-squares/QP (dense $\mathcal{O}(N(p{+}q)^{3})$ vs. sparse/structured factorizations) is based on [RAO2001, RAO2003] and the MPC text [RAWLINGS2017]. The real-time iteration viewpoint that justifies the sparse/condensed complexity discussion is due to [DiehlBockSchloeder2005]. Parallel time is reported as an order-of-growth proxy for the critical path using the work–span (Brent) bound,
@@
with $W$ the total work, $D$ the span (depth), and $P$ the number of processors [Brent1974].
@@
The proposed HNN estimators are a compelling middle ground: they inherit the parallel efficiency of dense linear–algebra workloads, enforces constraints by construction, mitigates disturbance bias via compensation neurons, and comes with explicit GUUB guarantees. Against the three baselines commonly used in practice [ASKARI2016, REICHBACH2016, MURAKAMI1998, ZHU1999, HASSAN2009, UHLICH2010, WANG2018, LUO2012, JOUKOV2015, HRUSTIC2020, SONG2022, RAO2001, RAO2001, RAO2003, RAWLINGS2017, ZAVALA2008], it offers a competitive compromise between performance and online complexity, especially when implemented on parallel hardware.